# Explicit monotone stable super-time-stepping methods for finite time singularities

### v1

@@ Abstract
We explore a novel way to numerically resolve the scaling behavior of finite-time singularities in solutions of nonlinear parabolic PDEs. The Runge–Kutta–Legendre (RKL) and Runge–Kutta–Gegenbauer (RKG) super-time-stepping methods were originally developed for nonlinear complex physics problems with diffusion. These are multi-stage single step second-order, forward-in-time methods with no implicit solves. The advantage is that the timestep size for stability scales with stage number $s$ as $\mathcal{O}(s^{2})$. Many interesting nonlinear PDEs have finite-time singularities, and the presence of diffusion often limits one to using implicit or semi-implicit timestep methods for stability constraints. Finite-time singularities are particularly challenging due to the large range of scales that one desires to resolve, often with adaptive spatial grids and adaptive timesteps. Here we show two examples of nonlinear PDEs for which the self-similar singularity structure has time and space scales that are resolvable using the RKL and RKG methods, without forcing even smaller timesteps. Compared to commonly-used implicit numerical methods, we achieve significantly smaller run time while maintaining comparable accuracy. We also prove numerical monotonicity for both the RKL and RKG methods under their linear stability conditions for the constant coefficient heat equation, in the case of infinite domain and periodic boundary condition, leading to a theoretical guarantee of the superiority of the RKL and RKG methods over traditional super-time-stepping methods, such as the Runge-Kutta-Chebyshev (RKC) and the orthogonal Runge-Kutta-Chebyshev (ROCK) methods.
@@ 1 Introduction
Many numerical approaches have been designed to resolve the theoretical scaling analysis of finite time singularities. These include both adaptive mesh refinement with (semi-)implicit schemes analysis [5, 13, 14, 8, 9] and schemes based on dynamic rescaling in similarity variables [4, 27, 28, 40]. Semi-implicit methods may be slower than explicit methods since large systems of implicit equations need to be solved for each update. Whereas explicit methods can save the computational cost of solving large systems of equations, they often have timestep restrictions for stability.
@@ 2.1 Semilinear Heat Equation
The semilinear heat equation 1 is a well-known PDE with finite-time singularity [4, 10, 32]; more precisely, with non-negative initial condition $u_{0}(x)$, there exists some $T>0$ such that $\underset{t\to T}{\lim}\|u(x,t)\|_{\infty}=\infty$. The asymptotic behavior of the semilinear heat equation has been studied extensively; a generic stable formula was first proposed by M. Berger and R. Kohn [4] to be
@@
which shows that the maximum value of $u(x,t)$ is linear with respect to the time remaining before the blow-up in the $\log$-$\log$ scale, with slope $-\frac{1}{p-1}$.
In order to perform an accurate numerical simulation, it is essential to understand and resolve the intrinsic scaling in the solution of the PDE. Consider the half-width of the solution $x_{\text{half}}\equiv x_{\text{half}}(t)$ defined by $u(x_{\text{half}}(t),t)=\frac{1}{2}\|u(\cdot,t)\|_{L^{\infty}}$. By definition and the previous observation,
$\frac{1}{2}\Big{[}(T-t)(p-1)\Big{]}^{-\frac{1}{p-1}}\sim\frac{1}{2}\|u(\cdot,t)\|_{L^{\infty}}\sim(T-t)^{\frac{-1}{p-1}}\Big{[}(p-1)+\frac{(p-1)^{2}}{4p}\frac{x_{\text{half}}^{2}}{(T-t)|\log(T-t)|}\Big{]}^{\frac{-1}{p-1}}.$
Direct calculation gives $2^{p-1}\sim 1+(p-1)\frac{x_{\text{half}}^{2}}{4p(T-t)|\log(T-t)|}$, so that
$x_{\text{half}}^{2}\sim\frac{4p(T-t)|\log(T-t)|(2^{p-1}-1)}{p-1}.$
Taking the logarithm on both sides, we have that $\log(x_{\text{half}}^{2})\sim\log(T-t)+\log(|\log(T-t)|4p(2^{p-1}-1))-\log(p-1)$; since the first term dominates over the second term when $T-t$ is small enough, we choose to neglect the second term in the asymptotic behavior when testing the numerical result, and therefore get the estimation
(5) $\log(x_{\text{half}}^{2})\sim\log(T-t)-\log(p-1)\sim(1-p)\log\|u(\cdot,t)\|_{L^{\infty}}-2\log(p-1)$
so that indeed $x^{2}\sim T-t$ and there is a linear relationship between $\log(x_{\text{half}}^{2})$ and $\log\|u(\cdot,t)\|_{L^{\infty}}$, with slope $-\frac{1}{p-1}$ with respect to the independent variable $\log(x_{\text{half}}^{2})$. In the numerical simulation near the blow-up time, one must resolve the scaling $dt\sim dx^{2}$ entailed by the PDE itself with logarithmic correction; this indicates that we should use small number of substages $s$ in the stabilized Runge-Kutta methods near the blow-up time. In order to resolve the structure of the singularity we need to choose $dt\sim\mathcal{O}(dx^{2})$, which allows us to use a smaller $s$ near the blow-up, resulting in more computational efficiency.
@@ 2.2 Surface Diffusion Equation
In contrast to the blow-up behavior of the semilinear heat equation, the similarity solution 8 has a pinchoff at $z=0$ in finite-time, $\underset{t\to T}{\lim}\min r(z,t)=0$, with non-negative initial condition symmetric with respect to the origin. Therefore, a natural characteristic scaling within the PDE itself is $z^{4}\sim(T-t)$, and any numerical simulation needs to resolve the correct scaling $dt\sim dx^{4}$ near the pinchoff time to obtain the correct solution. Another result in [5] shows that there exists a countable family of self-similar symmetric pinchoff solutions $\overline{R}_{i}(\eta)$, $i\in\mathbb{N}_{0}$, each with a different far-field cone angle $\overline{R}_{i}(\eta)\sim\pm c_{i}\eta$ as $\eta\to\pm\infty$, such that $c_{i}>c_{i+1}$ for all $i\in\mathbb{N}_{0}$. The half-cone angles of the self-similar solutions are given by $\varphi_{i}=\arctan(c_{i})$, where the physically relevant solution is the one with $i=0$. Therefore, we expect our numerical simulation to capture the similarity solution with the correct cone angle. Finally, we also want to simulate the rate of change of the similarity solution. By a calculation in [5], we know that $\overline{R}_{0}(0)^{4}/4=0.060575684$, therefore $\Big{|}\frac{d}{dt}\overline{R}_{0}(0)\Big{|}=0.060575684/\overline{R}_{0}(0)^{3}$; taking the $\log$-$\log$ scale, we clearly observe that the rate of change of the minimum value scales like the negative of the minimum value cubed.
@@
Figure 1: Adaptive mesh refinement strategy similar to the one used in [8]. The above three plots show the process of two mesh refinements around the origin. Each vertical |||| represents a grid point, and the shorter |||| at each level represents the new grid points added.
@@ 4 Experiment results
We now present the computational results. We compare the accuracy and run time of our explicit super-time-stepping methods with the backward Euler implicit method. Various initial grid sizes are used to guarantee the robustness of the RKL and RKG methods. We present representative results here and more detailed experiment results can be found in the supplementary documents LABEL:app:experiments.
@@ 4.1 Semilinear Heat Equation
We consider the semilinear heat equation 1 for the cases $p=2$ and $p=3$. The spatial domain is $[-1,1]$ with symmetry about the origin. In both cases, the Dirichlet boundary condition $u(\pm 1,t)=0$ is used. The initial condition is $u(x,0)=u_{0}(x)=\frac{10}{1-0.5\cos(\pi x)}-\frac{20}{3}$ for all cases. Periodic boundary condition or other non-negative even initial conditions can also work equally well for our numerical method, as long as the initial maximum value $\|u_{0}\|_{L^{\infty}}$ is not overly small or large. The discretization and the entire algorithm are discussed in detail in Section 3.2. The initial values of $dt$ and $dx$ satisfy $dt=\frac{1}{8}dx$, and we use initial $dx=\frac{1}{128},\frac{1}{256},\frac{1}{512}$ as test cases. The spatial grid sizes are chosen to be powers of 2, so that in their digital representation is exact. The initial stage number $s$ is 200, guaranteeing in all cases that $dt/dx^{2}$ are in the stability region of the super-time-stepping method.
(a) RKL2 $\log$-$\log$ relationship between $T-t$ and maximum value
(b) RKG2 $\log$-$\log$ relationship between $T-t$ and maximum value
Figure 2: Log-log relationship between the $T-t$ and maximum value $\|u(\cdot,t)\|_{\infty}$, where $T$ is the theoretical blow-up time, with $p=3$. The calculated values fit perfectly with the predicted line, which has the expected slope $-\frac{1}{p-1}$.
Fig. 2 shows a $\log$-$\log$ relation between the time to blow-up and the maximum value; the predicted value comes from Equation 4. The plot is a straight line with slope $-\frac{1}{p-1}$ and horizontal intercept $-\frac{1}{p-1}\log(p-1)$. Since the exact blow-up time $T$ is unknown, we use the final time of our numerical simulation to approximate the theoretical blowing up time.
(a) RKL2 $\log$-$\log$ relationship between half-width and maximum value
(b) RKG2 $\log$-$\log$ relationship between half-width and maximum value
Figure 3: Log-log relationship between the squared half-width and the maximum of the solution, with $p=3$. The value $x_{\text{half}}$ satisfies that $u(x_{\text{half}},t)=\frac{1}{2}\|u(\cdot,t)\|_{\infty}$. The calculated values fit well with the predicted line, which has expected slope $-\frac{1}{p-1}$.
Fig. 3 shows the $\log$-$\log$ relationship between numerically calculated half-width and the maximum value of the solution; the slope of the plots are as expected. The half-width provides a natural measurement of the spatial scaling of the blow-up. We define it to be the spatial width of the solution where it reaches exactly half of the maximum value. Our goal is to track the similarity variable $\frac{x^{2}}{T-t}$. Here we expect the square of the half-width to be proportional to the maximum value of the numerical solution, with slope $\frac{1}{1-p}$ according to 5. Finally, we note that the formula for the similarity solution includes a log correction in half-width. However, as noted in Fig. 3, the log correction is infinitesimal over the large range of scales shown here.
@@ 4.2 Surface Diffusion Equation
Simulations of the surface diffusion equation 2 are conducted. First we verify the cone angle near the pinchoff. As mentioned in [5], the cone angle of the first analytic similarity solution is $46.0444^{\circ}$. This corresponds to the first derivative of the Eulerian variable solution near the pinchoff point and also to the far field derivative of the similarity solution, yielding a slope of approximately 1.037 in absolute value. The spatial domain is chosen to be $[-2\pi,2\pi]$ and periodic boundary conditions are used. The initial condition is $u(x,0)=-1.8(\frac{1}{4}\cos(\frac{x}{2})+\frac{1}{4})+1.2$, so that it is symmetric about the origin with moderately small minimum value. We test our numerical method with initial $dx=\frac{4\pi}{128},\frac{4\pi}{256},\frac{4\pi}{512}$ and initial $dt=10^{-5}$ for all the cases. Two and one initial mesh refinements are applied the first two cases respectively before the simulation for more accurate simulation near the singularity.
@@
Figure 9: Log-log relationship between the absolute value of time derivative of the minimum value and the minimum value of $r$, surface diffusion equation. The time derivative is numerically calculated using the standard forward difference. According to the theoretical prediction, after taking the logarithm the absolute value of the time derivative has slope $-3$.
dx=$\frac{4\pi}{128}$ dx=$\frac{4\pi}{256}$ dx=$\frac{4\pi}{512}$
@@
Table 2: Run time comparison between implicit Euler’s method and the super-time-stepping explicit method for the surface diffusion equation. For each initial $dx$, the fastest method is highlighted. We stop the simulation when the solution blows up to $2\times 10^{-4}$. All the super-time-stepping method simulations require significantly less computational time than the classical implicit Euler’s method.
@@ 5.1 Preliminary Definitions and Setup
where $\alpha\in\mathbb{R}_{+}$, $u\in\mathbb{R}^{\mathbb{Z}}$, $M:\mathbb{R}^{\mathbb{Z}}\to\mathbb{R}^{\mathbb{Z}}$ is defined by $[Mu]_{i}:=\frac{[u]_{i+1}-2[u]_{i}+[u]_{i-1}}{dx^{2}}$, $[u]_{i}$ is the $i$-th entry of $u\in\mathbb{R}^{\mathbb{Z}}$, $i\in\mathbb{Z}$. We denote the CFL coefficient by $c:=\alpha\frac{dt}{dx^{2}}$, where $dt$ is the super timestep size. Notice that all the results below also work for the periodic and homogeneous Neumann boundary condition cases, since the periodic boundary condition problem can be equivalently seen as a boundaryless problem with repeated values, and the homogeneous Neumann boundary condition problem can be seen as a periodic boundary condition problem with each period to be the original domain reflected and doubled.
@@ 5.2 Runge–Kutta–Legendre Methods Monotonicity
The difficulty of the proof lies in controlling the Legendre polynomial and in the repetitive application of the central difference operator $M$ on the grid values. It is natural to start with an induction attempt given the recurrence relation of the Legendre polynomial as in the derivation of the RKL methods. However, for the RKL methods, there is barely any relationship between the $s$-stage method and the $(s+1)$-stage method. For example, for the RKL1 method, $P_{s+1}\Big{(}1+\frac{2}{(s+1)^{2}+(s+1)}x\Big{)}$ and $P_{s}\Big{(}1+\frac{2}{s^{2}+s}x\Big{)}$ cannot be directly linked by the Legendre polynomial recurrence relation given the different arguments in the Legendre polynomial. Another approach might be to write an RKL method as the composition of Euler’s method with different timestep lengths [34], and try to prove the monotonicity of each Euler’s method. Yet each of the internal Euler’s step might not even be stable, thus leading to no hope of its monotone stability, which in general is a stronger property than the numerical stability. We also note that Shu-Osher [43] represented the classical RK methods in a similar formulation to the RKL and RKG method. They proposed a simple criterion for the RK method to be total-variation diminishing (TVD) [29]. However, with more careful observation, it can be seen that the Shu-Osher representation cannot be directly applied to the RKL or the RKG methods, since it requires the coefficients before the intermediate steps $Y_{j}$ in 16 to be all non-negative, a property that is neither satisfied by the RKL nor the RKG method; moreover, the TVD property is a weaker condition than our desired monotone stability [29].

### v2

@@ Abstract
We explore a novel way to numerically resolve the scaling behavior of finite-time singularities in solutions of nonlinear parabolic PDEs. The Runge–Kutta–Legendre (RKL) and Runge–Kutta–Gegenbauer (RKG) super-time-stepping methods were originally developed for nonlinear complex physics problems with diffusion. These are multi-stage single step second-order, forward-in-time methods with no implicit solves. The advantage is that the timestep size for stability scales with stage number $s$ as $\mathcal{O}(s^{2})$. Many interesting nonlinear PDEs have finite-time singularities, and the presence of diffusion often limits one to using implicit or semi-implicit timestep methods for stability constraints. Finite-time singularities are particularly challenging due to the large range of scales that one desires to resolve, often with adaptive spatial grids and adaptive timesteps. Here we show two examples of nonlinear PDEs for which the self-similar singularity structure has time and space scales that are resolvable using the RKL and RKG methods, without forcing even smaller timesteps. Compared to commonly-used implicit numerical methods, we achieve significantly smaller run time while maintaining comparable accuracy. We also prove numerical monotonicity for both the RKL and RKG methods under their linear stability conditions for the constant coefficient heat equation, in the case of infinite domain and periodic boundary condition, leading to a theoretical guarantee of the superiority of the RKL and RKG methods over traditional super-time-stepping methods, such as the Runge-Kutta-Chebyshev (RKC) and the orthogonal Runge-Kutta-Chebyshev (ROCK) methods. Code can be found at https://github.com/ZT220501/SRK-Singularity.
@@ 1 Introduction
Many numerical approaches have been designed to resolve the theoretical scaling analysis of finite time singularities. These include both adaptive mesh refinement with (semi-)implicit schemes analysis [5, 13, 14, 8, 9] and schemes based on dynamic rescaling in similarity variables [4, 27, 28, 40, 22]. Semi-implicit methods may be slower than explicit methods since large systems of implicit equations need to be solved for each update. Whereas explicit methods can save the computational cost of solving large systems of equations, they often have timestep restrictions for stability.
@@ 2.1 Semilinear Heat Equation
The semilinear heat equation 1 is a well-known PDE with finite-time singularity [4, 10, 32]; more precisely, with non-negative initial condition $u_{0}(x)$, there exists some $T>0$ such that $\underset{t\to T}{\lim}\|u(x,t)\|_{\infty}=\infty$. The asymptotic behavior of the semilinear heat equation has been studied extensively; a generic stable self-similar formulation was first proposed by M. Berger and R. Kohn [4] to be
@@
which shows that the maximum value of $u(x,t)$ is linear with respect to the time remaining before the blow-up in the $\log$-$\log$ scale, with slope $\frac{1}{p-1}$.
In order to perform an accurate numerical simulation, it is essential to understand and resolve the intrinsic scaling in the solution of the PDE. The numerical simulation near the blow-up time must resolve the scaling $dt\sim dx^{2}$ entailed by the PDE itself with logarithmic correction; this indicates that we should use small number of substages $s$ in the stabilized Runge-Kutta methods near the blow-up time. In order to resolve the structure of the singularity we need to choose $dt\sim\mathcal{O}(dx^{2})$, which allows us to use a smaller $s$ near the blow-up, resulting in more computational efficiency.
@@ 2.2 Surface Diffusion Equation
In contrast to the blow-up behavior of the semilinear heat equation, the similarity solution 7 has a pinchoff at $z=0$ in finite-time, $\underset{t\to T}{\lim}\min r(z,t)=0$, with non-negative initial condition symmetric with respect to the origin. Therefore, a natural characteristic scaling within the PDE itself is $z^{4}\sim(T-t)$, and any numerical simulation needs to resolve the correct scaling $dt\sim dx^{4}$ near the pinchoff time to obtain the correct solution. Another result in [5] shows that there exists a countable family of self-similar symmetric pinchoff solutions $\overline{R}_{i}(\eta)$, $i\in\mathbb{N}_{0}$, each with a different far-field cone angle $\overline{R}_{i}(\eta)\sim\pm c_{i}\eta$ as $\eta\to\pm\infty$, such that $c_{i}>c_{i+1}$ for all $i\in\mathbb{N}_{0}$. The half-cone angles of the self-similar solutions are given by $\varphi_{i}=\arctan(c_{i})$, where the physically relevant solution is the one with $i=0$. Therefore, we expect our numerical simulation to capture the similarity solution with the correct cone angle. Finally, we also want to simulate the rate of change of the similarity solution. By a calculation in [5], we know that $\overline{R}_{0}(0)^{4}/4=0.060575684$, therefore $\Big|\frac{d}{dt}\overline{R}_{0}(0)\Big|=0.060575684/\overline{R}_{0}(0)^{3}$; taking the $\log$-$\log$ scale, we clearly observe that the rate of change of the minimum value scales like the negative of the minimum value cubed.
@@
Figure 1: Adaptive mesh refinement strategy similar to the one used in [8]. The above three plots show the process of two mesh refinements around the origin. Each vertical $|$ represents a grid point, and the shorter $|$ at each level represents the new grid points added.
@@ 4 Experiment results
We now present the computational results. We compare the accuracy and run time of our explicit super-time-stepping methods with the backward Euler implicit method. Various initial grid sizes are used to guarantee the robustness of the RKL and RKG methods. We present representative results here and more detailed experiment results can be found in the supplementary documents.
@@ 4.1 Semilinear Heat Equation
We consider the semilinear heat equation 1 for the cases $p=2$ and $p=3$. The spatial domain is [1,1][-1,1] with symmetry about the origin. In both cases, the Dirichlet boundary condition $u(\pm 1,t)=0$ is used. The initial condition is $u(x,0)=u_{0}(x)=\frac{10}{1-0.5\cos(\pi x)}-\frac{20}{3}$ for all cases. Periodic boundary condition or other non-negative even initial conditions can also work equally well for our numerical method, as long as the initial maximum value $\|u_{0}\|_{L^{\infty}}$ is not overly small or large. The discretization and the entire algorithm are discussed in detail in Section 3.2. The initial values of $dt$ and $dx$ satisfy $dt=\frac{1}{8}dx$, and we use initial $dx=\frac{1}{128},\frac{1}{256},\frac{1}{512}$ as test cases. The spatial grid sizes are chosen to be powers of 2, so that in their digital representation is exact. The initial stage number $s$ is 200, guaranteeing in all cases that $dt/dx^{2}$ are in the stability region of the super-time-stepping method.
(a) RKL2 $\log$-$\log$ relationship between TtT-t and maximum value
(b) RKG2 $\log$-$\log$ relationship between TtT-t and maximum value
Figure 2: Log-log relationship between the TtT-t and maximum value $\|u(\cdot,t)\|_{\infty}$, where $T$ is the theoretical blow-up time, with $p=3$. The calculated values fit perfectly with the predicted line, which has the expected slope $\frac{1}{p-1}$.
Fig. 2 shows a $\log$-$\log$ relation between the time to blow-up and the maximum value; the predicted value comes from Equation 4. The plot is a straight line with slope $\frac{1}{p-1}$ and horizontal intercept $\frac{1}{p-1}\log(p-1)$. Since the exact blow-up time $T$ is unknown, we use the final time of our numerical simulation to approximate the theoretical blowing up time.
(a) RKL2 similarity variable scaling
(b) RKG2 similarity variable scaling
Figure 3: Log-log relationship between the squared half-width and $(T-t)|\log(T-t)|$, with $p=3$. The value $x_{\text{half}}$ satisfies that $u(x_{\text{half}},t)=\frac{1}{2}\|u(\cdot,t)\|_{\infty}$. The calculated values fit well with the predicted line, which has expected slope around 11, indicating that the numerical solution accurately captures the scaling of the semilinear heat equation with the log correction.
Fig. 3 shows the $\log$-$\log$ relationship between the square of the numerically calculated half-width and the time variable quantity $(T-t)|\log(T-t)|$. The half-width provides a natural measurement of the spatial scaling of the blow-up. Our goal is to track the similarity variable $\frac{x^{2}}{(T-t)|\log(T-t)|}$. According the the similarity profile in Equation 3, we expect the plot to indicate a linear relationship with slope $1$. The results in Fig. 3 confirms our expectation, indicating that our numerical simulations accurately captures the similarity profile.
@@ 4.2 Surface Diffusion Equation
Simulations of the surface diffusion equation 2 are conducted. First we verify the cone angle near the pinchoff. As mentioned in [5], the cone angle of the first analytic similarity solution is $46.0444^{\circ}$. This corresponds to the first derivative of the Eulerian variable solution near the pinchoff point and also to the far field derivative of the similarity solution, yielding a slope of approximately 1.037 in absolute value. The spatial domain is chosen to be $2\pi,2\pi]$ and periodic boundary conditions are used. The initial condition is $u(x,0)=-1.8(\frac{1}{4}\cos(\frac{x}{2})+\frac{1}{4})+1.2$, so that it is symmetric about the origin with moderately small minimum value. We test our numerical method with initial $dx=\frac{4\pi}{128},\frac{4\pi}{256},\frac{4\pi}{512}$ and initial $dt=10^{-5}$ for all the cases. Two and one initial mesh refinements are applied the first two cases respectively before the simulation for more accurate simulation near the singularity.
@@
Figure 9: Log-log relationship between the absolute value of time derivative of the minimum value and the minimum value of $r$, surface diffusion equation. The time derivative is numerically calculated using the standard forward difference. According to the theoretical prediction, after taking the logarithm the absolute value of the time derivative has slope 3-3.
[…]
@@
Table 2: Run time comparison between implicit Euler’s method and the super-time-stepping explicit method for the surface diffusion equation. For each initial dxdx, the fastest method is highlighted. We stop the simulation when the solution blows up to $2\times 10^{-4}$. All the super-time-stepping method simulations require significantly less computational time than the classical implicit Euler’s method.
@@ 5.1 Preliminary Definitions and Setup
where $\alpha\in\mathbb{R}_{+}$, $u\in\mathbb{R}^{\mathbb{Z}}$, $M:\mathbb{R}^{\mathbb{Z}}\to\mathbb{R}^{\mathbb{Z}}$ is defined by $Mu]_{i}:=\frac{[u]_{i+1}-2[u]_{i}+[u]_{i-1}}{dx^{2}}$, $u]_{i}$ is the $i$-th entry of $u\in\mathbb{R}^{\mathbb{Z}}$, $i\in\mathbb{Z}$. We denote the CFL coefficient by $c:=\alpha\frac{dt}{dx^{2}}$, where $dt$ is the super timestep size. Notice that all the results below also work for the periodic and homogeneous Neumann boundary condition cases, since the periodic boundary condition problem can be equivalently seen as a boundaryless problem with repeated values, and the homogeneous Neumann boundary condition problem can be seen as a periodic boundary condition problem with each period to be the original domain reflected and doubled.
@@ 5.2 Runge–Kutta–Legendre Methods Monotonicity
The difficulty of the proof lies in controlling the Legendre polynomial and in the repetitive application of the central difference operator $M$ on the grid values. It is natural to start with an induction attempt given the recurrence relation of the Legendre polynomial as in the derivation of the RKL methods. However, for the RKL methods, there is barely any relationship between the $s$-stage method and the $(s+1)$-stage method. For example, for the RKL1 method, $P_{s+1}\Big(1+\frac{2}{(s+1)^{2}+(s+1)}x\Big)$ and $P_{s}\Big(1+\frac{2}{s^{2}+s}x\Big)$ cannot be directly linked by the Legendre polynomial recurrence relation given the different arguments in the Legendre polynomial. Another approach might be to write an RKL method as the composition of Euler’s method with different timestep lengths [34], and try to prove the monotonicity of each Euler’s method. Yet each of the internal Euler’s step might not even be stable, thus leading to no hope of its monotone stability, which in general is a stronger property than the numerical stability. We also note that Shu-Osher [43] represented the classical RK methods in a similar formulation to the RKL and RKG method. They proposed a simple criterion for the RK method to be total-variation diminishing (TVD) [29]. However, with more careful observation, it can be seen that the Shu-Osher representation cannot be directly applied to the RKL or the RKG methods, since it requires the coefficients before the intermediate steps $Y_{j}$ in 15 to be all non-negative, a property that is neither satisfied by the RKL nor the RKG method; moreover, the TVD property is a weaker condition than our desired monotone stability [29].